Differentiate the function.f(x) = log[sub]9[/sub](xe[sup]x[/sup])
1
Expert's answer
2011-02-25T06:00:51-0500
f'(x) = [(xex)ln9]-1* (xex)' = ex(1 +x)* [(xex)ln9]-1= (1+x)/(x ln9)